### v1

@@ 3.9 Nested
3.10 Continuity
Another concern that has been raised [18] is that with some methodologies, the interval may be a discontinuous function of the choice of confidence level, or the $p$ value may be a discontinuous function of the null hypothesis [28]. A slight change in confidence level could lead to a large change in the interval, or a slight change in null hypothesis could lead to a large change in p value. Since, for any given observation, the pdf is a continuous function of $\theta$, the confidence interval might be expected to be continuous in the confidence level. Two models which are essentially identical could result in large differences in $p$ value if continuity is violated. Thus, continuity of the interval with respect to confidence level, or of the $p$ value as a function of null hypothesis, is desirable.
3.11 Gives sensible $p$ values
A good confidence interval will also yield sensible $p$ values. For example, the $p$-value should be a monotonic function of the observation for observations on either side of $H_{0}$. That is, it should be a bimonotonic function of $n$, for any $H_{0}$ not at the boundary of parameter space. This property is related to other properties we have listed.
@@ 3.12 Discussion
It has been shown [22] that it is not generally possible to simultaneously achieve properties of continuous intervals, optimal lengths, and nesting. [^6] Violation of nesting from inverting a hypothesis test can also happen with continuous distributions [29]. We must decide on what is most important.
Since we usually discuss intervals in the context of a given choice of confidence level, it might be argued that nesting and continuity do not appear as a practical matter. However, we sometimes do look at different confidence level intervals for a given observation, and then we could encounter issues. The issue also becomes crisper when we look at $p$ values. We return to this in Section 10.
@@
Often, we know on physical grounds that $\theta$ cannot be negative, even though the mathematically allowed parameter space in Eq. 5 extends to $\theta=-b$. This need not be a concern until one makes statements about the true value of $\theta$, which we are not doing; we are only describing a measurement. Nevertheless, sometimes people are uncomfortable with the idea that the description could include “unphysical” values in the parameter space. Of course, if this happens for some $n$, we can just cut the interval off to exclude the unphysical values. This is perfectly fair, because it does not invalidate the probability statement in Eq. 6. However, we recommend against doing this truncation, because it actually makes the descriptive statement more difficult to understand, that is, it obscures the intuitive connection with the sampled $n$. Nonetheless, physicists have invented special techniques incorporating the background in a manner that avoids dealing with unphysical regions, and we examine two popular such approaches in sections 8.1 and 8.2.
@@ 4 One-sided intervals
The discussion of upper and lower limits is very similar; we’ll consider the concrete case of an upper limit here. If we want to construct an upper limit, as our confidence interval, there really isn’t much ambiguity for a frequentist limit. For a Poisson distribution with mean $\mu=\theta+b$ we look for sets of the form
@@
$P(N\in S_{\alpha}(\mu)|\mu)\geq 1-\alpha.$ (28)
@@
$P(N\geq n_{\rm min}|\mu)=\sum_{n=n_{\rm min}}^{\infty}f(n;\mu)\geq 1-\alpha.$ (29)
@@
$C_{\alpha}(n)=\{\mu|n\in S_{\alpha}(\mu)\}=(0,\mu=\mu_{\rm max}),$ (30)
where $\mu_{\rm max}$ is determined as the smallest $\mu$ such that
@@
Besides being smallest, we can check whether our standard upper limit satisfies the other desirable properties we have listed as well. It is by construction an exact interval, satisfying Eq. 6. It is also connected. It is not guaranteed to contain the maximum likelihood estimator. For example, if we observe a large $n$, the maximum likelihood estimator is $\hat{\mu}=n$, but if $1-\alpha$ is small the upper limit may be less than $n$. However, as long as $1-\alpha\geq 1/2$, the upper limit is guaranteed to include $n$, the maximum likelihood estimator, as may readily be demonstrated numerically. We may demonstrate optimal coverage, i.e., Eq. 19. Among possible upper limits, as we have noted in the construction, it gives intervals of shortest length. The upper limit would satisfy the asymptotic requirement if we rephrased the requirement for a one-sided interval. Our upper limit is ordered in the sense that, if $n_{2}>n_{1}$, then the corresponding upper limits have the property $u_{2}>u_{1}$. The property of “symmetry” is meaningless for a one-sided interval. Our upper limits are strictly nested, that is, if $1-\alpha_{2}>1-\alpha_{1}$ then $u_{2}>u_{1}$ for any given observation. The upper limit is a continuous function of $\alpha$.
@@
$\begin{split}\hbox{Prob}(X<2\mu)&=\int_{0}^{2\mu}\frac{x^{n-1}e^{-x/2}}{\Gamma(n)2^{n}}dx\\ \hbox{Prob}(N\geq n)&=1-\sum_{k=0}^{u}{n-1}\frac{\mu^{k}e^{-\mu}}{k!}.\end{split}$ (32)
@@
We note that our upper limits on $\theta$ (when background is present) could end up being negative, even if it is known that $\theta\geq 0$ on physical grounds. This is not an issue, the limits are performing as designed (that is, as frequency intervals). If you nonetheless find this concerning, I urge you to re-examine your goal in computing a limit. The fact that you care about consistency with physical values of $\theta$, that is with “truth”, strongly suggests a goal of interpretation, not description. As discussed in the Introduction, “truth” is not the domain of frequency statistics. If the goal is to make a statement concerning how large the true value of $\theta$ might be, this is interpretation. For example, when designing a new experiment to search for a rare process, one starts with some idea for how large the true value might be. This informs the design of the new effort, including decisions around whether it is worth pursuing. Resort to Bayesian methods.
In fact, frequency upper limits are not typically especially interesting as descriptions of a measurement. Our two-sided intervals are much more informative. However, the upper limit we have described provides the often useful one-sided $p$ value discussed at the end of section 2.2.
@@ 5 Conventional methods
For all of the methods discussed in this section, the intervals in $\mu$ and $\theta$ are simply related. If $\ell,u)$ is a confidence interval for $\mu$ then $\ell-b,u-b)$ is the interval for $\theta$ at the same confidence level. Thus, we need only discuss one parameter, which we’ll take to be $\mu$.
@@ 5.1 The “equal-tailed” or “fiducial” interval
Putting together two one-sided intervals (upper and lower), we may obtain two-sided intervals, and this is the earliest approach to defining a two-sided confidence interval for the Poisson distribution. The classic Poisson confidence interval is attributed to Garwood [19] [^7] Sometimes rediscovered in specific fields, e.g., [30].. This commonly used exact approach to confidence intervals emphasizes symmetry (in probability) about the observed counts. The idea is to combine a lower limit with CL $1-\alpha/2$ with an upper limit with CL $1-\alpha/2$ to obtain a central interval with CL $1-\alpha$.
Given $\mu$ let the set $S^{+}_{\alpha}(\mu)$ be the set of all observations such that $\mu$ is contained in the $1-\alpha/2$ upper confidence limit. This is given by
$S^{+}_{\alpha}(\mu)=\{n|\hbox{Prob}(N\leq n|\mu)\geq 1-\alpha/2\}.$ (36)
Likewise, let the set $S^{-}_{\alpha}(\mu)$ be the set of all observations such that $\mu$ is contained in the $1-\alpha/2$ lower confidence limit. This is given by
$S^{-}_{\alpha}(\mu)=\{n|\hbox{Prob}(N\geq n|\mu)\geq 1-\alpha/2\}.$ (37)
Then for any given observation $N=n$, a two-sided $1-\alpha$ confidence interval for $\mu$ is given by
$C_{\alpha}(n)=\{\mu|n\in S^{-}_{\alpha}(\mu)\hbox{ and }n\in S^{+}_{\alpha}(\mu)\}.$ (38)
Thus, for given $N=n$, the confidence interval is $(\ell,u)$, where
$\begin{split}\sum_{k=n}^{\infty}\frac{\ell^{k}}{k!}e^{-\ell}&=\alpha/2\\ \sum_{k=0}^{n}\frac{u^{k}}{k!}e^{-u}&=\alpha/2.\end{split}$ (39)
Following the discussion in Sec. 4, this may be computed according to:
$\left(\ell(n),u(n)\right)=\left(F^{-1}(\alpha/2;2n)/2,F^{-1}(1-\alpha/2;2(n+1))/2\right),$ (40)
@@
Figure 2: Performance of Garwood intervals. Left: coverage probability as a function of $\mu$; Middle: 68% confidence interval as a function of observed counts. The plus symbols show the maximum likelihood estimator for the mean; Right: size of the confidence interval as a function of observed counts; Right: size of the 68% confidence interval as a function of observed counts. The plus symbols show the values of $2\sqrt{n}$.
@@
The Garwood interval is exact, connected, strictly nested, ordered, and continuous in the confidence level. It is symmetric in probability by construction. It also has the desired asymptotic property and leads to bimonotonic $p$ values. Thus, it meets many of our desired criteria.
@@
It is possible to reduce the overcoverage by relaxing the requirement that each tail have probability $\geq 1-\alpha/2$. That is, we can directly invert a two-sided test, referred to [18] as “strictly two-sided intervals”. Thus, several two-sided intervals (discussed below) with better coverage properties have been devised, abandoning the symmetry property. On the other hand, Thulin and Zwanzig [18] remark that the Garwood interval is optimal (i.e., shortest) among strictly nested intervals.
@@ 5.2 Smallest size $S_{\alpha}(\mu)$ by ordering on probability
The idea of Eqs. 41 and 42 is that, for given $\mu$, we find those values of $n$ (there are either one or two such values) for which the pdf is maximal. These values are added to $S_{\alpha}(\mu)$. Then we look for the next highest values of $f$ and add the corresponding values of $n$ to $S_{\alpha}(\mu)$. This is continued until we get a set with probability at least $1-\alpha$.
@@
Given a sampled value, $n$, we look through sets $S_{\alpha}(\mu)$ for those which contain $n$. Because of the monotonicity of $n_{\ell}$ and $n_{h}$ with respect to $\mu$, the sets that contain $n$ will include a continuous range of $\mu$. Thus, our confidence set is an interval, $C_{\alpha}(n)=(\ell,u)$, where $n_{\ell}=\max_{n}S_{\alpha}(\ell)$ and $n_{h}=\min_{n}S_{\alpha}(u)$. The analog of this method was proposed by Sterne [24] for the binomial distribution, and the method has been discussed for the Poisson distribution by Crow and Gardner [20]. We refer to this interval as “Sterne”.
@@
Crow and Gardner [20] proposed a variation on this idea that preserves the minimum size of $S_{\alpha}(\mu)$ but also minimizes the upper bound on the confidence interval among all such smallest sets instead of strictly ordering on probability. We refer to this interval as “Crow&Gardner”. It may be noted that this modification destroys the correspondence of the Crow&Gardner intervals with a test-inversion. Instead, effectively a different test is defined for each $\alpha$. This suggests that deriving $p$ values from these intervals may be ill-defined. We’ll discuss this further in Section 10.
@@
Schilling and Holladay find, with their alternative to the Casella and Robert length criterion, (section 3.5) that the Crow&Gardner intervals [20] are optimal on this criterion, but also have endpoints that are not strictly increasing in $n$. They argue for a modified procedure, designated “Modified Crow-Gardner (MCG)”, to obtain stricly increasing endpopints while keeping small the resulting violations of length optimality: Start with limits $\ell_{\rm MCG}(n)=\ell_{\rm CG}(n)$ for all $n$, where “CG” indicates the Crow-Gardner intervals. Starting with $n=0$, whenever $\ell_{\rm MCG}(n+1)<ml(n)\equiv\ell_{\rm MCG}(n)+\min\left[0.01\ell_{\rm MCG}(n),0.1)\right]$, make the increase $\ell_{\rm MCG}(n+1)=ml(n)$. The corresponding coincidental endpoint, in which $u_{\rm CG}(m)=\ell_{\rm CG}(n+1)$ for some $m$, is simultaneously increased.
@@ 5.3 Blaker
(Blaker) Define function $\gamma(\mu,n)\equiv\min\left[P(N\geq n|\mu),P(N\leq n|\mu)\right]$ and “acceptability function” $\alpha(\mu;n)\equiv P\left[\gamma(\mu,N)\leq\gamma(\mu,n)|\mu\right]$. Then
@@
      If $P(N\geq n|\mu)<P(N\leq n|\mu)$, then
      $\alpha(\mu;n)=P(N\geq n|\mu)+P(N\leq n^{*}|\mu),$ (44)
      where $n^{*}$ is the largest $m$ such that $P(N\leq m|\mu)\leq P(N\geq n|\mu)$
@@
      If $P(N\geq n|\mu)=P(N\leq n|\mu)$, then
@@
      If $P(N\geq n|\mu)>P(N\leq n|\mu)$, then
      $\alpha(\mu;n)=P(N\leq n|\mu)+P(N\geq n^{**}|\mu),$ (46)
      where $n^{**}$ is the smallest $k$ such that $P(N\geq k|\mu)\leq P(N\leq n|\mu)$.
@@
When we look at continuity and nestedness, we find that the Blaker intervals are nested, though not strictly, as a function of confidence level, but exhibit discontinuous behavior, Fig. 8a. The Crow&Gardner intervals are continuous, but not nested, Fig. 8b. Comparing these figures, we may decide that nesting is a more important property to require – at least the intervals never get smaller with increasing confidence level.
@@
Figure 8: Confidence intervals as a function of confidence level $1-\alpha$, for observation $n=2$

### v2

@@ 4.9 Nested
4.10 Continuity and monotonicity
Another concern that has been raised [1] is that with some methodologies, the interval may be a discontinuous function of the choice of confidence level, or the $p$-value may be a discontinuous function of the null hypothesis [48]. A slight change in confidence level could lead to a large change in the interval, or a slight change in null hypothesis could lead to a large change in $p$-value. Since, for any given observation, the pdf is a continuous function of $\theta$, the confidence interval might be expected to be continuous in the confidence level. Two models which are essentially identical could result in large differences in $p$-value if continuity is violated. Thus, continuity of the interval with respect to confidence level, or of the $p$-value as a function of null hypothesis, is desirable.
We also consider it important for the upper and lower interval bounds to be monotonic functions of the confidence level, $1-\alpha$. Otherwise we could have the counterintuitive situation in which the upper bound decreases or the lower bound increases when the CL increases. It also seems desirable for the interval length to monotonically increase as $n$ increases, but we do not consider this to be as important, at least if exceptions are not too glaring.
4.11 Gives sensible $p$-values
A good confidence interval will also yield sensible $p$-values. For example, the $p$-value should be a monotonic function of the observation for observations on either side of $H_{0}$. That is, it should be a bimonotonic function of $n$, for any $H_{0}$ not at the boundary of parameter space. This property is related to other properties we have listed.
@@ 4.12 Discussion
It has been shown [4] that, for discrete distributions such as Poisson and binomial, it is not generally possible to simultaneously achieve properties of continuous intervals, optimal lengths, and nesting. [^8] Violation of nesting from inverting a hypothesis test can also happen with continuous distributions [7]. We must decide on what is most important. Since we usually discuss intervals in the context of a given choice of confidence level, it might be argued that nesting and continuity do not appear as a practical matter. However, we sometimes do look at different confidence level intervals for a given observation, and then we could encounter issues. The issue also becomes crisper when we look at $p$-values. We return to this in section 11.
@@
Often, we know on physical grounds that $\theta$ cannot be negative (or cannot be less than, or perhaps not be greater than some value) even though the mathematically allowed parameter space in Eq. 1 extends to $\theta=-b$. This need not be a concern until one makes statements about the true value of $\theta$, which we are not doing; we are only describing a measurement. Nevertheless, sometimes people are uncomfortable with the idea that the description could include “unphysical” values in the parameter space. Of course, if this happens for some $n$, we can just cut the interval off to exclude the unphysical values. This is mathematically fair, because it does not invalidate the probability statement in Eq. 6. However, we recommend against doing this truncation, because it actually makes the descriptive statement more difficult to understand, that is, it obscures the intuitive connection with the sampled $n$. Nonetheless, physicists have invented special techniques incorporating the background in a manner that avoids dealing with unphysical regions, and we examine two popular such approaches in sections 9.1 and 9.2.
@@ 5 One-sided intervals
The discussion of upper and lower limits is very similar; we will consider the concrete case of an upper limit here. If we want to construct an upper limit, as our confidence interval, there really is not much ambiguity for a frequentist limit. For a Poisson distribution with mean $\mu=\theta+b$ we look for sets of the form
@@
$\hbox{Prob}(N\in S_{\alpha}(\mu)|\mu)\geq 1-\alpha.$ (29)
@@
$\hbox{Prob}(N\geq n_{\rm min}|\mu)=\sum_{n=n_{\rm min}}^{\infty}f(n;\mu)\geq 1-\alpha.$ (30)
@@
$C_{\alpha}(n)=\{\mu|n\in S_{\alpha}(\mu)\}=[0,\mu=\mu_{\rm max}],$ (31)
where $\mu_{\rm max}$ is determined as the largest $\mu$ such that
@@
Besides being smallest, we can check whether our standard upper limit satisfies the other desirable properties we have listed as well. It is by construction an exact interval, satisfying Eq. 6. It is also connected. It is not guaranteed to contain the maximum likelihood estimator. For example, if we observe a large $n$, the maximum likelihood estimator is $\hat{\mu}=n$, but if $1-\alpha$ is small the upper limit may be less than $n$. However, as long as $1-\alpha\geq 1/2$, the upper limit is guaranteed to include $n$, the maximum likelihood estimator, as may readily be demonstrated numerically. Since we usually concern ourselves with $1-\alpha\geq 1/2$, this feature may be considered to be acceptable. We may demonstrate optimal coverage, i.e., Eq. 20. Among possible upper limits, as we have noted in the construction, it gives intervals of shortest length. The upper limit would satisfy the asymptotic requirement if we rephrased the requirement for a one-sided interval. Our upper limit is ordered in the sense that, if $n_{2}>n_{1}$, then the corresponding upper limits have the property $u_{2}>u_{1}$. The property of “symmetry” is meaningless for a one-sided interval. Our upper limits are strictly nested, that is, if $1-\alpha_{2}>1-\alpha_{1}$ then $u_{2}>u_{1}$ for any given observation. The upper limit is a continuous function of $\alpha$.
@@
$\begin{split}\hbox{Prob}(X<2\mu)&=\int_{0}^{2\mu}\frac{x^{n-1}e^{-x/2}}{\Gamma(n)2^{n}}dx\\ \hbox{Prob}(N\geq n)&=1-\sum_{k=0}^{n-1}\frac{\mu^{k}e^{-\mu}}{k!}.\end{split}$ (33)
@@
We note that our upper limits on $\theta$ (when background is present) could end up being negative (or less than some physical minimum for $\theta$), even if it is known that $\theta\geq 0$ (or some other value) on physical grounds. This is not an issue, the limits are performing as designed (that is, as frequency intervals). If you nonetheless find this concerning, I urge you to re-examine your goal in computing a limit. The fact that you care about consistency with physical values of $\theta$, that is with “truth”, strongly suggests a goal of interpretation, not description. As discussed in the Introduction and section 3, “truth” is not the domain of frequency statistics. If the goal is to make a statement concerning how large the true value of $\theta$ might be, this is interpretation. For example, when designing a new experiment to search for a rare process, one starts with some idea for how large the true value might be. This informs the design of the new effort, including decisions around whether it is worth pursuing. Resort to Bayesian methods.
In fact, frequency upper limits are not typically especially interesting as descriptions of a measurement. Our two-sided intervals are much more informative. James and Roos [18] also point this out, including the observation that if “only an upper limit is published, the result cannot be combined with those of other experiments in a straightforward way…” However, the upper limit we have described provides the often useful one-sided $p$-value discussed at the end of section 2.2.
@@ 6 Conventional methods
For all of the methods discussed in this section, the intervals in $\mu$ and $\theta$ are simply related. If $\ell,u]$ is a confidence interval for $\mu$ then $\ell-b,u-b]$ is the interval for $\theta$ at the same confidence level. Thus, we need only discuss one parameter, which we will take to be $\mu$. Again, $b$ is a known quantity – we are not thinking, for example, about a difference of two Poisson RVs.
@@ 6.1 The “equal-tailed” or “fiducial” interval
Putting together two one-sided intervals (upper and lower), we may obtain two-sided intervals, and this is the earliest approach to defining a two-sided confidence interval for the Poisson distribution. The classic Poisson confidence interval is attributed to Garwood [15] [^9] Sometimes rediscovered in specific fields, e.g., [46]., who derived it as a fiducial interval. It is readily checked that it satisfies equation 6 (e.g., [32]). This commonly used approach to confidence intervals emphasizes symmetry (in probability) about the observed counts. The idea is to combine a lower limit with CL $1-\alpha/2$ with an upper limit with CL $1-\alpha/2$ to obtain a central interval with CL $1-\alpha$.
As a fiducial interval, we start with Eq. 4 and find, for given $N=n$, the confidence interval $\ell(n),u(n)]$, where
$\begin{split}\hbox{Prob}[N\geq n|\mu=\ell(n)]&=\sum_{k=n}^{\infty}\frac{\ell^{k}}{k!}e^{-\ell}=\alpha/2\\ \hbox{Prob}[N\leq n|\mu=u(n)]&=\sum_{k=0}^{n}\frac{u^{k}}{k!}e^{-u}=\alpha/2.\end{split}$ (37)
If $n=0$, we set $\ell=0$. Following the discussion in Sec. 5, this may be computed according to:
$\left[\ell(n),u(n)\right]=\left[F^{-1}(\alpha/2;2n)/2,F^{-1}(1-\alpha/2;2(n+1))/2\right],$ (38)
@@
Figure 2: Performance of Garwood intervals. Left: coverage probability as a function of $\mu$; Middle: 68% confidence interval as a function of observed counts. The plus symbols show the maximum likelihood estimator for the mean; Right: size of the 68% confidence interval as a function of observed counts. The plus symbols show the values of $2\sqrt{n}$.
@@
The Garwood interval is exact, connected, strictly nested, ordered, and continuous in the confidence level. It is symmetric in probability by construction. It also has the desired asymptotic property and leads to bimonotonic $p$-values. Thus, it meets many of our desired criteria.
@@
It is possible to reduce the overcoverage by relaxing the requirement that each tail have probability $\geq 1-\alpha/2$. That is, we can directly invert a two-sided test, referred to in [1] as “strictly two-sided intervals”. Thus, several two-sided intervals (discussed below) with better coverage properties have been devised, abandoning the symmetry property. On the other hand, Thulin and Zwanzig [1] remark that the Garwood interval is optimal (i.e., shortest) among strictly nested intervals.
@@ 6.2 Smallest size $S_{\alpha}(\mu)$ by ordering on probability
The idea of Eqs. 39 and 40 is that, for given $\mu$, we find those values of $n$ (there are either one or two such values) for which the pdf is maximal. These values are added to $S_{\alpha}(\mu)$. Then we look for the next highest values of $f$ and add the corresponding values of $n$ to $S_{\alpha}(\mu)$. This is continued until we get a set with probability at least $1-\alpha$.
@@
Given a sampled value, $n$, we look through sets $S_{\alpha}(\mu)$ for those which contain $n$. Because of the monotonicity of $n_{\ell}$ and $n_{h}$ with respect to $\mu$, the sets that contain $n$ will include a continuous range of $\mu$. Thus, our confidence set is an interval, $C_{\alpha}(n)=[\ell,u]$, where $n_{\ell}=\max_{n}S_{\alpha}(\ell)$ and $n_{h}=\min_{n}S_{\alpha}(u)$. The analog of this method was proposed by Sterne [41] for the binomial distribution, and the method has been discussed for the Poisson distribution by Crow and Gardner [10]. We refer to this interval as “Sterne”.
@@
Crow and Gardner [10] proposed a variation on this idea that preserves the minimum size of $S_{\alpha}(\mu)$ but also minimizes the upper bound on the confidence interval among all such smallest sets instead of strictly ordering on probability. We refer to this interval as “Crow&Gardner”. It may be noted that this modification destroys the correspondence of the Crow&Gardner intervals with a test-inversion. Instead, effectively a different test is defined for each $\alpha$. This suggests that deriving $p$-values from these intervals may be ill-defined. We will discuss this further in section 11.
@@
Schilling and Holladay [37] find, with their alternative to the Casella and Robert length criterion, (section 4.5) that the Crow&Gardner intervals [10] are optimal on this criterion, but also have endpoints that are not strictly increasing in $n$. They argue for a modified procedure, designated “Modified Crow-Gardner (MCG)”, to obtain strictly increasing endpoints while keeping small the resulting violations of length optimality: Start with limits $\ell_{\rm MCG}(n)=\ell_{\rm CG}(n)$ for all $n$, where “CG” indicates the Crow-Gardner intervals. Starting with $n=0$, whenever $\ell_{\rm MCG}(n+1)<ml(n)\equiv\ell_{\rm MCG}(n)+\min\left[0.01\ell_{\rm MCG}(n),0.1)\right]$, make the increase $\ell_{\rm MCG}(n+1)=ml(n)$. The corresponding coincidental endpoint, in which $u_{\rm CG}(m)=\ell_{\rm CG}(n+1)$ for some $m$, is simultaneously increased.
@@ 6.3 Blaker
(Blaker) Define function $\gamma(\mu,n)\equiv\min\left[\hbox{Prob}(N\geq n|\mu),\hbox{Prob}(N\leq n|\mu)\right]$ and “acceptability function” $\alpha(\mu;n)\equiv\hbox{Prob}\left[\gamma(\mu,N)\leq\gamma(\mu,n)|\mu\right]$. Then
@@
      If $\hbox{Prob}(N\geq n|\mu)<\hbox{Prob}(N\leq n|\mu)$, then
      $\alpha(\mu;n)=\hbox{Prob}(N\geq n|\mu)+\hbox{Prob}(N\leq n^{*}|\mu),$ (42)
      where $n^{*}$ is the largest $m$ such that $\hbox{Prob}(N\leq m|\mu)\leq\hbox{Prob}(N\geq n|\mu)$
@@
      If $\hbox{Prob}(N\geq n|\mu)=\hbox{Prob}(N\leq n|\mu)$, then
@@
      If $\hbox{Prob}(N\geq n|\mu)>\hbox{Prob}(N\leq n|\mu)$, then
      $\alpha(\mu;n)=\hbox{Prob}(N\leq n|\mu)+\hbox{Prob}(N\geq n^{**}|\mu),$ (44)
      where $n^{**}$ is the smallest $k$ such that $\hbox{Prob}(N\geq k|\mu)\leq\hbox{Prob}(N\leq n|\mu)$.
@@
When we look at continuity and nestedness, we find that the Blaker intervals are nested, though not strictly, as a function of confidence level, but exhibit discontinuous behavior, Fig. 8a. The Crow&Gardner intervals are continuous, but neither nested nor monotonic with CL, Fig. 8b. Comparing these figures, we may decide that sacrificing nesting to obtain continuity is not a good trade-off if it comes at the expense of monotonicity.
@@
Figure 8: Confidence intervals as a function of confidence level $1-\alpha$, for observation $n=2$. The curves show the upper and lower bounds in the intervals.